Obtain the Riccati equation associated with the equation y"+w^2 y=0 and hence find its solution?
1
Expert's answer
2019-09-27T09:22:11-0400

First, let's consider some of the equalities associated with taking derivative of composite functions

"\frac{d}{dx}y^2=2y⋅y′"

"\frac{d}{dx}y'^2=2y'⋅y''"

Then, multiplying both sides by "y'"

"y 
′′
 +ω^2
 y=0|| \times 2y 
′"

we obtain

"2y \n\u2032\n \u22c5y \n\u2032\u2032\n +\u03c9^\n2\n \u22c5(2y \n\u2032\n \u22c5y)=0"

"\\frac{d}{dx}\n\n\u200b\t\n (y'^2\n )+ \n\\frac{d}{dx}\n\n\u200b\t\n (\u03c9^2\n \u22c5y^2\n )=0"

Therefore, applying the sum rule

"\\frac{d}{dx}\n\n\u200b\t\n (y'^2\n +\u03c9^2\n \u22c5y^2\n )=0"

"y'^2\n +\u03c9^2\n \u22c5y^2=C^2"

Resolving this equation we get

"y'=\pm\sqrt{C^2-ω^2
 ⋅y^2}"

"\frac{dy}{\sqrt{C^2-ω^2
 ⋅y^2}}=\pm dx"

"\int \frac{dy}{\sqrt{C^2-ω^2
 ⋅y^2}}=\pm x+C_1"

We introduce a substitution

"u=\\frac{\u03c9\u22c5y}{C}\n\u200b"

"du=\frac{ω⋅dy}{C}"

"\int \frac{Cdu}{ωC\sqrt{1-u^2}}=\pm x+C_1"

"\frac{1}{ω}\arcsin(u)=\pm x+C_1"

"\arcsin(u)=ω(\pm x+C_1)"

"u=\sin(ω((\pm x+C_1)))"

"\frac{ω⋅y}{C}=\sin(ω((\pm x+C_1))"

"y=\frac{C}{ω}\sin(ω((\pm x+C_1))"